# Dynamic Moiré-like pattern in non-Hermitian Wannier-Stark ladder system

### v1

@@ Abstract
I Introduction
Wannier-Stark (WS) ladders are a theoretical model in solid-state physics that describe the quantization of electronic states in a crystal under the influence of a constant electric field Bloch1929 ; Wannier1959 ; Wannier1960 ; Glueck2002 ; Waschke1993 . It is closely related to Bloch oscillations (BOs) due to the equidistant spectrum. In the past decade, such a phenomenon has attracted much attention in cold-atoms physics and photonics due to applications in interferometric measurements and as a method for manipulating localized wave packets Breid2006 ; Breid2007 ; Dreisow2009 ; Kling2010 ; Ploetz2011 . It can be simulated using artificial quantum systems, such as superconducting circuits Song2024 . On the other hand, WS ladders are an essential theoretical tool for understanding and exploring particle transport phenomena in quantum systems and have potential applications in superconducting quantum computing and quantum simulation. In the previous study, encompassing both theoretical and experimental aspects, only a linear external field was considered because it is central to the formation of an equidistant spectrum.
@@ I Introduction
One of the unique features of a non-Hermitian system is the violation of conservation law of the Dirac probability, based on which, the complex potential and non-Hermitian hopping strength are employed to describe open systems phenomenologically J2004 . Furthermore, unconventional propagation of light associated with the gain/loss has been demonstrated by engineering effective non-Hermitian Hamiltonians in optical systems Guo2009 ; Makris2008 ; Musslimani2008 ; Klaiman2008 ; Longhi2009 ; ElGanainy2007 ; Zheng2010 .
Exceptional points (EPs), as an exclusive feature of a non-Hermitian system, are degeneracies of non-Hermitian operators Berry2004 ; Heiss2012 ; Miri2019 ; Zhang2020 . The corresponding eigenstates coalesce into one state resulting to the incompleteness of Hilbert space. The peculiar features around EP have sparked tremendous attention to the classical and quantum photonic systems Doppler2016 ; Xu2016 ; Assawaworrarit2017 ; Wiersig2014 ; Wiersig2016 ; Hodaei2017 ; Chen2017 .Notably, a coalescing state has an exclusive feature. Both theoretical and experimental works not only give an insight into the dynamical property of the non-Hermitian Hamiltonian but also provide a platform to implement the novel optical phenomenon.
@@
In general, Moiré patterns emerge from the superposition of two periodic structures that have either slightly different periods or different orientations. These patterns have been realized in materials Yankowitz2012 ; Ponomarenko2013 ; Dean2013 ; Hunt2013 ; Woods2014 ; Nakajima2016 ; Lohse2015 . Recently, the influence of Moiré patterns in physical systems has attracted considerable interest Ponomarenko2013 ; Dean2013 ; Hunt2013 ; Gorbachev2014 ; Song2015 ; Jung2014 , particularly in non-Hermitian systems and the dynamics of such systems Yang2018 ; Wang2020 . The Moiré-like patterns presented in this work are based on a somewhat different mechanism. They mainly arise from the quasi periodicity resulting from the non-linear potential.
@@
This paper is organized as follows. In Section II, the non-Hermitian SSH model under a position-dependent potential is introduced. We explore the symmetries of the Hamiltonian with a linear potential and investigate the possible structures of energy levels based on these symmetries. In Section III and Section IV, we employ the non-Hermitian Floquet method to calculate the phase diagram. In Section V, the dynamical behavior for the case with a linear potential is studied. We also propose a model with non-linear potentials and demonstrate the Moiré-like pattern in the dynamics through numerical simulations. A summary is provided in the Section VI.
Figure 1: Schematic illustrations of (a) the non-Hermitian SSH Hamiltonian with a tilted field in Eq. (1), which consists of adjacent real hopping strength $J$ and imaginary hopping strength $i\beta$, and (b-d) its consponding energy level structures for three typical cases. When the on-site tilted potential has a uniform slope $\omega$, the Hamiltonian obeys the three symmetries defined by Eqs. (2), (5), and (10), respectively. These constrain the energy spectrum partition the energy spectrum into three distinct categories: (b) two real ladders with an energy shift, forming a symmetric spectrum with respect to zero energy; (c) a single coalescing ladder with level spacing $2\omega$, (d) two complex conjugate pair ladders. The black solid lines represent the real part of the energy levels, while the pink and blue blocks represent the positive imaginary part and negative imaginary part of the energy levels, respectively.
II Hamiltonian and EP Wannier Stark ladder
@@
We consider a non-Hermitian variant of SSH chain Su1980 ; Takayama1980 ; Roth1987 ; Jackiw1976 ; Heeger1988 ; Ryu2002 ; Hasan2010 ; Qi2011 imposed by a position dependent field with the Hamiltonian
$H$ $=$ $\sum\limits_{j=-\infty}^{+\infty}[J(|2j\rangle\langle 2j+1|+\,\mathrm{h.c.})+i\beta(|2j\rangle\langle 2j-1|$ (1)
$+\,\mathrm{h.c.})]+\sum\limits_{l=-\infty}^{+\infty}V_{l}|l\rangle\langle l|]\,,$
@@
where $|l\rangle$ denotes a site state describing the Wannier state localized on the $l$th period of the potential. Here, $J$ and $i\beta$ are the tunneling strength, and $V_{l}$ a on-site potential. In this paper, we take $J$ and $\beta>0$ for the sake of simplicity. The system is schematically illustrated in Fig. 1(a). In previous work, it has been shown that the energy levels obeys a simple structures in its real and imaginary parts when taking the linear potential, $V_{l}=\omega(l+\frac{1}{2})$ Zhang2024 . In this work, we consider the case involving a non-linear potential $V_{l}$, which will be demonstrated to give rise to a dynamic Moiré-like pattern.
@@
and $\mathcal{T}i\mathcal{T}^{-1}=-i$. According to non-Hermitian quantum mechanics Scholtz1992 , a $\mathcal{PT}$-symmetric Hamiltonian, as a pseudo-Hermitian system, is characterized by energy levels that appear either real or in complex conjugate pairs. The eigenstates corresponding to complex conjugate energy levels are interconnected by the $\mathcal{PT}$ operation. Introducing the translation operator $T_{2}$ defined as
@@
III Non-Hermitian Floquet method
In this section, we investigate the energy level structure within the framework of the non-Hermitian Floquet method, which represents an extension of the conventional Floquet method Maksimov2015 to the presnt non-Hermitian Hamiltonian.
@@ III Non-Hermitian Floquet method
$2\omega(l+\frac{1}{4})\phi_{l}^{A}-J\phi_{l}^{B}-i\beta\phi_{l-1}^{B}$ $=$ $E\phi_{l}^{A},$ (16)
$2\omega(l+\frac{3}{4})\phi_{l}^{B}-J\phi_{l}^{A}-i\beta\phi_{l+1}^{A}$ $=$ $E\phi_{l}^{B}.$
@@
Figure 2: (a) The phase diagram of the Hamiltonian, as depicted in Eq. (1), illustrating the different phases of the system. The green and blue areas correspond to the parameter ranges where the real non-degenerate energy levels in Fig. 1(b) and the complex conjugate pairs of energy levels in Fig. 1(d) occur, respectively. The red circles represent the distribution of the EP boundary between phases, while the red solid line is the result predicted by the Eq. (39). (b) The same phase diagram as (a) is presented with an additional specific line, and four typical points are labeled. The dotted line represents the parameters with $\beta=0.46J$, varying $J/\omega$ from $9.8$ (point A) to $16.2$ (point D). This dotted line goes through two distinct phases, indicating the transitions from one state to another periodically. The dynamic behaviors for the systems with parameters corresponding to points C and B are presented in Figs. 3(a) and 3(b), respectively. This line provides a way to demonstrate different dynamic behaviors along a chain with the non-linear potential given by Eq. (47), exhibiting Moiré-like patterns as shown in Fig. 4.
IV Phase diagram
IV.1 Phase diagram
@@
For a system with a broad range of parameters $(J,\beta,\omega)$, the types of Stark ladders in an infinite chain can be ascertained through numerical simulations employing the Floquet method. In Fig. 2(a) we plot the phase diagram in the $J/\omega-\beta/\omega$ plane.
IV.2 The distribution of EPs under the weak-field approximation
@@
$h_{\mathrm{I}}$ $=$ $e^{-ih_{0}k}h^{\prime}e^{ih_{0}k}$ (35)
$=$ $-\left(\begin{array}[]{cc}0&Je^{ik/2}+i\beta e^{-ik/2}\\ Je^{-ik/2}+i\beta e^{ik/2}&0\end{array}\right).$
@@
$\mathbf{\Phi}(\pi/\omega)=-\exp i(\varphi_{\mathrm{D}}+\varphi_{\mathrm{B}})\sigma_{z}\mathbf{\Phi}(0),$
@@
Figure 3: The profiles of the evolved states for the initial state given by Eq. (45) under the systems with parameters corresponding to points C and B in Fig. 2(b). 3D plots of (a) $P_{n}(t)$ and (b) $P_{n}/P$, as defined in Eqs. (43) and (46), are obtained by exact diagonalization for finite system with $N=100$. We observe that the evolved state (a) is a superposition of two Bloch oscillations with a period $T=\pi/\omega$, and (b) represents the amplified Bloch oscillations as expected. The logarithm of the total Dirac probability, $\text{In}P(t)$, indicates that difference between two kinds of dynamic behaviors is evident.
V Dynamics Under Linear and Nonlinear Potentials
@@
$|\Phi(0)\rangle=\sum_{l,\sigma=\pm}c_{l}^{\sigma}|\psi_{l}^{\sigma}\rangle,$ (41)
@@
where $\left\{|\psi_{l}^{\pm}\rangle\right\}$ is normalized with Dirac inner product by taking $\left|\langle\psi_{l}^{\pm}|\psi_{l}^{\pm}\rangle\right|^{2}=1$. (i) In the case with $E_{0}^{+}=\left|E_{0}^{+}\right|$, it is clear that the evolved state is the superposition of two Bloch oscillations with a period of $\pi/\omega$. We are interested in the Dirac probability of the evolved state, which is defined by
@@
$|\Phi(t)\rangle\approx e^{E_{0}^{+}t}\sum_{l}c_{l}^{+}e^{-i2l\mathcal{\omega}t}|\psi_{l}^{+}\rangle,$ (44)
@@
$P_{n}(t)=\left|\left\langle n\right|\Phi(t)\rangle\right|^{2}.$ (46)
@@
Figure 4: The profile of the evolved state, given the initial state in Eq. (48), is computed for the system with $\beta=0.46J$ ($J=1$) and a non-linear potential as described by Eq. (47). The 3D plot of $P_{n}(t)$, as defined in Eq. (46), shows that the evolving state exhibits evident Moiré-like patterns. The paremeters are $T=\pi\tau=13\pi$,$\ \gamma=1.6\times 10^{-3}$, and $N=4000$.
@@
$V_{l}=\frac{1}{\gamma l+\tau}(l+\frac{1}{2}),$ (47)
where the constants $\gamma$ and $\tau$ satisfy $\gamma\ll 1/J\ll\tau$. It allows mapping the parameter line in Fig. 2(b) onto an SSH chain, varying the local $J/\omega$ from $9.8$ to $16.2$ along the chain. We take the initial state being distributed on each site with the equal probability
@@
We plot the probability distributions $P_{n}(t)$ in Fig. 4. It can be observed that the evolving state exhibits apparent Moiré-like patterns. Before ending this paper, it is worth briefly mentioning that the Moiré-like pattern previously discussed can be implemented through an array of evanescently coupled optical waveguides Makris2008 ; Musslimani2008 ; Klaiman2008 ; Longhi2009 ; ElGanainy2007 . In this context, the time evolution of state $\left|\Phi(t)\right\rangle$ mapped into the spatial evolution of the modal amplitudes of light waves, along the array axis.
VI Summary

### v2

@@ Abstract
keywords:
Non-Hermitian quantum mechanics , Exceptional points ladder , Non-linear field , Moiré-like patterns
PACS:
03.65.-w , 11.30.Er , 71.10.Fd
Email address: ]songtc@nankai.edu.cn
1 Introduction
Wannier-Stark (WS) ladders are a theoretical model in solid-state physics that describe the quantization of electronic states in a crystal under the influence of a constant electric field [1, 2, 3, 4, 5]. It is closely related to Bloch oscillations (BOs) due to the equidistant spectrum. In the past decade, such a phenomenon has attracted much attention in cold-atoms physics and photonics due to applications in interferometric measurements and as a method for manipulating localized wave packets [6, 7, 8, 9, 10]. It can be simulated using artificial quantum systems, such as superconducting circuits [11]. On the other hand, WS ladders are an essential theoretical tool for understanding and exploring particle transport phenomena in quantum systems and have potential applications in superconducting quantum computing and quantum simulation. In the previous study, encompassing both theoretical and experimental aspects, only a linear external field was considered because it is central to the formation of an equidistant spectrum.
@@ 1 Introduction
One of the unique features of a non-Hermitian system is the violation of conservation law of the Dirac probability, based on which, the complex potential and non-Hermitian hopping strength are employed to describe open systems phenomenologically [12]. Furthermore, unconventional propagation of light associated with the gain/loss has been demonstrated by engineering effective non-Hermitian Hamiltonians in optical systems [13, 14, 15, 16, 17, 18, 19].
Exceptional points (EPs), as an exclusive feature of a non-Hermitian system, are degeneracies of non-Hermitian operators [20, 21, 22, 23]. The corresponding eigenstates coalesce into one state resulting to the incompleteness of Hilbert space. The peculiar features around EP have sparked tremendous attention to the classical and quantum photonic systems [24, 25, 26, 27, 28, 29, 30]. Notably, a coalescing state has an exclusive feature. Both theoretical and experimental works not only give an insight into the dynamical property of the non-Hermitian Hamiltonian but also provide a platform to implement the novel optical phenomenon.
@@
In general, Moiré patterns emerge from the superposition of two periodic structures that have either slightly different periods or different orientations. These patterns have been realized in materials [31, 32, 33, 34, 35, 36, 37]. Recently, the influence of Moiré patterns in physical systems has attracted considerable interest [32, 33, 34, 38, 39, 40], particularly in non-Hermitian systems and the dynamics of such systems [41, 42]. In comparison with existing work on Moiré patterns, our work has the following characteristics: (i) The Moiré-like patterns presented in this work are based on a somewhat different mechanism. They mainly arise from the quasi-periodicity resulting from a single periodic structure with a non-linear potential, rather than from two periodic structures. (ii) The Moiré-pattern can be detected in the dynamics rather than in static properties. (iii) This is the first time we apply the Floquet method to a non-Hermitian system.
@@
This paper is organized as follows. In Section 2, the non-Hermitian SSH model under a position-dependent potential is introduced. We explore the symmetries of the Hamiltonian with a linear potential and investigate the possible structures of energy levels based on these symmetries. In Section 3 and Section 4, we employ the non-Hermitian Floquet method to calculate the phase diagram. In Section 5, the dynamical behavior for the case with a linear potential is studied. We also propose a model with non-linear potentials and demonstrate the Moiré-like pattern in the dynamics through numerical simulations. A summary is provided in the Section 6.
2 Hamiltonian and EP Wannier Stark ladder
@@
We consider a non-Hermitian variant of SSH chain [43, 44, 45, 46, 47, 48, 49, 50] imposed by a position dependent field with the Hamiltonian
$H=\sum\limits_{j=-\infty}^{+\infty}[J(|2j\rangle\langle 2j+1|+\,\mathrm{h.c.})+i\beta(|2j\rangle\langle 2j-1|+\,\mathrm{h.c.})]+\sum\limits_{l=-\infty}^{+\infty}V_{l}|l\rangle\langle l|]\,,$ (1)
@@
where $|l\rangle$ denotes a site state describing the Wannier state localized on the $l$th period of the potential. Here, $J$ and $i\beta$ are the tunneling strength, and $V_{l}$ a on-site potential. In this paper, we take $J$ and $\beta>0$ for the sake of simplicity. The system is schematically illustrated in Fig. 1(a). In previous work, it has been shown that the energy levels obeys a simple structures in its real and imaginary parts when taking the linear potential, $V_{l}=\omega(l+\frac{1}{2})$ [51]. In this work, we consider the case involving a non-linear potential $V_{l}$, which will be demonstrated to give rise to a dynamic Moiré-like pattern.
Figure 1: Schematic illustrations of (a) the non-Hermitian SSH Hamiltonian with a tilted field in Eq. (1), which consists of adjacent real hopping strength $J$ and imaginary hopping strength $i\beta$ , and (b-d) its consponding energy level structures for three typical cases. When the on-site tilted potential has a uniform slope $\omega$, the Hamiltonian obeys the three symmetries defined by Eqs. (2), ( 5), and (10), respectively. These constrain the energy spectrum partition the energy spectrum into three distinct categories: (b) two real ladders with an energy shift, forming a symmetric spectrum with respect to zero energy; (c) a single coalescing ladder with level spacing $2\omega$, (d) two complex conjugate pair ladders. The black solid lines represent the real part of the energy levels, while the pink and blue blocks represent the positive imaginary part and negative imaginary part of the energy levels, respectively.
@@
and $\mathcal{T}i\mathcal{T}^{-1}=-i$. According to non-Hermitian quantum mechanics [52], a $\mathcal{PT}$-symmetric Hamiltonian, as a pseudo-Hermitian system, is characterized by energy levels that appear either real or in complex conjugate pairs. The eigenstates corresponding to complex conjugate energy levels are interconnected by the $\mathcal{PT}$ operation. Introducing the translation operator $T_{2}$ defined as
@@
3 Non-Hermitian Floquet method
In this section, we investigate the energy level structure within the framework of the non-Hermitian Floquet method, which represents an extension of the conventional Floquet method [53] to the presnt non-Hermitian Hamiltonian.
@@ 3 Non-Hermitian Floquet method
$2\omega(l+\frac{1}{4})\phi_{l}^{A}+J\phi_{l}^{B}+i\beta\phi_{l-1}^{B}$ $=$ $E\phi_{l}^{A},$ (16)
$2\omega(l+\frac{3}{4})\phi_{l}^{B}+J\phi_{l}^{A}+i\beta\phi_{l+1}^{A}$ $=$ $E\phi_{l}^{B}.$
@@
Figure 2: (a) The phase diagram of the Hamiltonian, as depicted in Eq. ( 1), illustrating the different phases of the system. The green and blue areas correspond to the parameter ranges where the real non-degenerate energy levels in Fig. 1(b) and the complex conjugate pairs of energy levels in Fig. 1(d) occur, respectively. The red circles represent the distribution of the EP boundary between phases, while the red solid line is the result predicted by the Eq. (38). (b) The same phase diagram as (a) is presented with an additional specific line, and four typical points are labeled. The dotted line represents the parameters with $\beta=0.46J$, varying $J/\omega$ from $8.6$ (point A) to $16.6$ (point D). This dotted line goes through two distinct phases, indicating the transitions from one state to another periodically. The dynamic behaviors for the systems with parameters corresponding to points C and B are presented in Figs. 3(a) and 3(b), respectively. This line corresponds to the parameter trajectory in Fig. 4(a), and provides a way to demonstrate different dynamic behaviors along a chain with the non-linear potential given by Eq. (46), exhibiting Moiré-like patterns as shown in Fig. 4(b).
4 Phase diagram
4.1 Phase diagram
@@
For a system with a broad range of parameters $(J,\beta,\omega)$, the types of Stark ladders in an infinite chain can be ascertained through numerical simulations employing the Floquet method. In Fig. 2(a) we plot the phase diagram in the $J/\omega-\beta/\omega$ plane. For a clearer demonstration, we have illustrated in Fig. 4(a) how the real and imaginary parts of the base energy level $E_{0}^{+}$ change along the trajectories of the two phases in the parameter space. The parameter trajectory is chosen as the line segment from point A to point D in Fig. 2(b), and it can be observed that $E_{0}^{+}$ oscillates between the two phases as the parameter varies.
4.2 The distribution of EPs under the weak-field approximation
@@
$h_{\mathrm{I}}=e^{-ih_{0}k}h^{\prime}e^{ih_{0}k}=-\left(\begin{array}[]{cc}0&Je^{ik/2}+i\beta e^{-ik/2}\\ Je^{-ik/2}+i\beta e^{ik/2}&0\end{array}\right).$ (33)
@@
$\mathbf{\Phi}(\pi/\omega)=-\exp i(\varphi_{\mathrm{D}}+\varphi_{\mathrm{B}})\sigma_{z}\mathbf{\Phi}(0),$ (35)
@@
Figure 3: The profiles of the evolved states for the initial state given by Eq. (44) under the systems with parameters corresponding to points C and B in Fig. 2(b). 3D plots of (a) $P_{l}(t)$ and (b) $P_{l}/P$, as defined in Eqs. (42) and (45), are obtained by exact diagonalization for finite system with $N=100$. We observe that the evolved state (a) is a superposition of two Bloch oscillations with a period $T=\pi/\omega$, and (b) represents the amplified Bloch oscillations as expected. The logarithm of the total Dirac probability, $\text{In}P(t)$, indicates that difference between two kinds of dynamic behaviors is evident.
5 Dynamics Under Linear and Nonlinear Potentials
@@
$|\Phi(0)\rangle=\sum_{n,\sigma=\pm}c_{n}^{\sigma}|\psi_{n}^{\sigma}\rangle,$ (40)
@@
where $\left\{|\psi_{n}^{\pm}\rangle\right\}$ is normalized with Dirac inner product by taking $\left|\langle\psi_{n}^{\pm}|\psi_{n}^{\pm}\rangle\right|^{2}=1$. (i) In the case with $E_{0}^{+}=\left|E_{0}^{+}\right|$, it is clear that the evolved state is the superposition of two Bloch oscillations with a period of $\pi/\omega$. We are interested in the Dirac probability of the evolved state, which is defined by
@@
$|\Phi(t)\rangle\approx e^{E_{0}^{+}t}\sum_{n}c_{n}^{+}e^{-i2n\mathcal{\omega}t}|\psi_{n}^{+}\rangle,$ (43)
@@
$P_{l}(t)=\left|\left\langle l\right|\Phi(t)\rangle\right|^{2}.$ (45)
@@
Figure 4: (a) The numerical results of the base energy level $E_{0}^{+}$ along the parameter trajectory of the line segment from point A to point D in Fig. 2(b). The green and blue blocks above represent the $\mathcal{PT}$-symmetric phase and the $\mathcal{PT}$-broken phase, respectively. it can be observed that $E_{0}^{+}$ oscillates between the two phases as the parameter varies. (b) The profile of the evolved state, given the initial state in Eq. (47), is computed for the system with $\beta=0.46J$ ($J=1$) and a non-linear potential as described by Eq. (46). The 3D plot of $P_{l}(t)$, as defined in Eq. (45), shows that the evolving state exhibits evident Moiré -like patterns. Obviously, the dark regions and bright strips correspond exactly to the $\mathcal{PT}$-symmetric phase and the $\mathcal{PT}$-broken phase areas in Figure (a), respectively. The paremeters are $T=\pi\tau=12.6\pi$,$\ \gamma=2\times 10^{-3}$, and $N=4000$.
@@
[…]
where the constants $\gamma$ and $\tau$ satisfy $\gamma\ll 1/J\ll\tau$ and $\omega=\text{{d}}V_{l}/\text{{d}}l={1}/(\gamma l+\tau)$ is position-dependent. It allows mapping the parameter line in Fig. 2(b) onto an SSH chain, varying the local $J/\omega$ from $8.6$ to $16.6$ along the chain. We take the initial state being distributed on each site with the equal probability
@@
We plot the probability distributions $P_{l}(t)$ in Fig. 4(b). The values of 1/$\omega$ in different lattice points of the SSH chain correspond to the horizontal axis of Fig. 4(a). It can be observed that the evolving state exhibits apparent Moiré-like patterns. The dark regions, characterized by a vanishing distribution, are interspersed between the bright strips, which exhibit a large distribution. From Fig. 4(a), it can be seen that these two regions correspond to the non-linear potential falling into two phases with distinct Bloch oscillations: the long-time dynamics will be dominated by the amplified Bloch oscillations, making the distribution in the regions with superimposed Bloch oscillations negligible.
Before ending this paper, we would like to discuss the scheme for experimentally detecting the Moiré-like pattern in photonic lattices. We consider a 1D array composed of $2N$ waveguides, which are assumed to be weakly coupled. According to the coupled-mode theory, light propagation can be described by a Schroinger-like equation, which is typically used to mimic the dynamics of a tight-binding system [54]. In the tight-binding approximation, the propagation along the $z$-direction of the optical wave in such a photonic lattice is described by coupled-mode equations. The propagation of the optical wave in this photonic lattice is described by the following coupled-mode equations
$\frac{\text{{d}}A_{l}(z)}{\text{{d}}z}=i\left(\kappa_{l,l-1}A_{l-1}+\kappa_{l,l+1}A_{l+1}\right),$ (48)
where $A_{l}(z)$ is the propagating amplitude in the $l$th waveguide, and $\kappa_{l,l\pm 1}$ is the coupling coefficient between adjacent waveguides. To mimic Bloch oscillations, the photonic lattice is bent to modify its spatial distribution of the optical potential. The curvature is seen as an inertial force acting on the optical wave. The corresponding equation is
$\frac{\text{{d}}A_{l}(z)}{\text{{d}}z}=i\left(\kappa_{l,l-1}A_{l-1}+\kappa_{l,l+1}A_{l+1}\right)+iF_{l}lA_{l},$ (49)
where $F_{l}$ is the inertial force acting on the optical wave in the $l$th waveguide. The key point related to the present topic is as follows. In a region consisting of a small number of waveguides, $F_{l}$ can be regarded as approximately $l$-independent. However, the curvature in the radial direction cannot be regarded as a constant, leading to a nonlinear potential when a large $N$ is considered. According to our theoretical analysis and the results of numerical simulations, the shift of $F_{l}$ across a single period of Moiré-like pattern is estimated as $\{[\mathrm{d}(F_{l}l)/\mathrm{d}t]|_{l=l_{0}+\Delta N}\}^{-1}-\{[\mathrm{d}(F_{l}l)/\mathrm{d}t]|_{l=l_{0}}\}^{-1}\approx 2$, with $\Delta N=1000$, from the expression of $F_{l}$, given by
$F_{l}=\frac{1}{0.002l}\ln(0.002l+12.6).$ (50)
Based on these parameters, the geometry of the sample, including the range of the bent radius, and the minimum number of waveguides, can be determined. Additionally, in principle, our findings could be simulated in these systems beyond photonic systems, such as the electric circuit system [55] and the superconducting circuit system [11].
6 Summary